OpenStudy (anonymous):

For a large cooking project, Arthur needs to put 18 cups of water into a large pot. He can transport a maximum of 4 cups from the sink in a single trip. The pot already contains 5 cups of water. How many trips are required for Arthur to fill the pot to 18 cups? A.3 B.4 C.5 D.6

OpenStudy (anonymous):

@twvogels

OpenStudy (whpalmer4):

how many cups does he need to transfer to fill up the pot to 18 cups?

OpenStudy (anonymous):

He needs 18 cups total, but already has 5 in the pot. So he needs 18-5=13 more cups. He can carry 4 cups per trip so he needs to make: \[\frac{13cups}{4\frac{cups}{trip}}=3.25\] He will need to make 4 trips. The last trip will not be a full 4 cups but he still needs to make it to get to 18. (B)

OpenStudy (anonymous):

thank u